OpenStudy (anonymous):

How many solutions will this system of equations have? x - 2y = 24 3x - 6y = 72 A) No solution B) Infinite solutions C) One solution D) Two solutions

OpenStudy (danjs):

hint, multiply the first equation through by 3

OpenStudy (danjs):

what do ya notice

OpenStudy (anonymous):

or solve for x in x - 2y = 24

OpenStudy (anonymous):

then substitute the x you got in that equation for the x in this equation: 3x−6y=72

OpenStudy (anonymous):

so if you have solved for x in the first equation I gave you, you should get: x=2y+24

OpenStudy (anonymous):

So the equation you should get after substituting would be: 3(2y+24)−6y=72

OpenStudy (anonymous):

then simplify

OpenStudy (anonymous):

to get 72=72

OpenStudy (anonymous):

then subtract 72 from each side to get 0=0, which means there are an infinite amount of solutions